JT:

What is the excluded value? a/b-2 a = 0 b = 2 b = - 2 none of the above

JT:

@NarWhalGod @563blackghost

JT:

@Shadow

Shadow:

Do you know what an excluded value is?

JT:

No tbh

NarWhalGod:

nope I dont

Shadow:

It's a number than makes the denominator zero. Since you cannot divide by 0 (since 0 cannot go into any number). So an excluded value is the number that a variable can't be, because it would make the denominator zero.

Shadow:

\[\frac{ t }{ o + 5 }\] The excluded value is o = -5, because if we input that in \[\frac{ t }{ -5 + 5 } \] \[\frac{ t }{ 0 } = undefined\]